OpenStudy (anonymous):

the diagonal of a square is 11cm. What is the perimeter of the square ?

OpenStudy (anonymous):

do you know the formula to calculate the perimeter of square its 4 * (length of side)

OpenStudy (zzr0ck3r):

you have a square with diagonal 11 use law of sines to get the length of the side sin(45)/x = sin(90)/11 x/11 = sqrt(2)/2 x = 11sqrt(2)/2 perimeter is 4x so 4*11sqrt(2)/2 = 22sqrt(2)

OpenStudy (zzr0ck3r):

does that make sense @1LiveToBreath

Directrix (directrix):

@zzr0ck3r You could use the 45-45-90 theorem in reverse. The diagonal of a square is the hypotenuse of a 45-45-90 triangle. So, the diagonal is square root of 2 times a side of the 45-45-90 triangle or the square in this case. One side of the square would be 11/sqrt(2) and then multiply that by 4. That is essentially what you did but this route may be a little faster. Remember this, too: The diagonal of a cube is sqrt(3) times one of its edges. That comes in handy.

OpenStudy (anonymous):

\[\sqrt{2}a=11\] where a is the side of square!

OpenStudy (anonymous):

\[a=\frac{ 11 }{ \sqrt{2} }\]\[perimeter=4*\frac{ 11 }{ \sqrt{2} }=\frac{ 44 }{ \sqrt{2} }=22\sqrt{2}\]